# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ 2.2 Eigenvalue problem
corresponding to a Fourier decomposition in the spatial and temporal domain with complex Fourier coefficients $\hat{f}(r)$, real azimuthal and vertical wavenumbers $m$ and $k$, respectively, and complex eigenfrequencies $\omega=\sigma+i\nu$. Due to the $\theta$-periodicity of the cylindrical equilibrium, $m$ is quantised to integer values, whereas the vertical wavenumber $k$ can take arbitrary values in this case. The set of eigenoscillations of this equilibrium is fully determined with the wavenumbers $m$ and $k$.
We calculated all MHD eigenspectra in this work using the Legolas [^1] https://legolas.science. The code version used in this work is open-source under the GPL-3.0 license and available on GitHub at https://github.com/n-claes/legolas/tree/v2.0.0. code developed by Claes et al. (2020), De Jonghe et al. (2022) and recently upgraded in terms of solver and memory performance (Claes & Keppens, 2023). It uses a finite element approach to solve the following generalised matrix eigenvalue problem obtained from the linearised MHD equations, where the Fourier coefficients become eigenfunctions:
@@
Here, $\mathbf{X}$ is an 8-component eigenvector before its Galerkin decomposition (dropping the hat on the Fourier coefficients), containing eigenfunctions $(\rho,v_{r},v_{\theta},v_{z},T,a_{r},a_{\theta},a_{z})$, with $\rho$ the density perturbation, $T$ the plasma temperature perturbation and $\mathbf{a}$ the magnetic vector potential, which fixes the perturbed magnetic field $\mathbf{B}=\nabla\times\mathbf{a}$ and enforces the monopole constraint $\nabla\cdot\mathbf{B}$ exactly. Legolas constructs the matrices A and B from the equilibrium variations, the chosen finite element base functions in the eigenmode representations, and the required physical effects like viscosity, resistivity, Hall MHD, etc. (De Jonghe et al., 2022). As this work is limited to ideal MHD, the matrices A and B are modified from the static, linear MHD case with only flow and gravity as additional effects. The matrix system is then solved using various available solvers depending on the requirements. In this work, we use the QR-invert algorithm to produce complete eigenspectra, and the Arnoldi shift-invert method to search for eigenvalues in the neighbourhood of some complex number (Claes & Keppens, 2023). The mathematical formulation of the eigenvalue problem is closed by fixing the boundary conditions at radius $1$ and $1+\delta$. By default, Legolas assumes perfectly conducting, solid walls as boundaries, requiring $v_{r}=0$ and $a_{\theta}=a_{z}=0$ at those locations. As explained in Goedbloed & Keppens (2022), these solid wall boundaries will affect the detailed variation of MRIs, while SARIs are fully agnostic of these boundary conditions, an important distinction between both mode categories.
To substantiate our Legolas results, we combine them in several figures with results obtained with the complementary ROC code, which employs a shooting method (Goedbloed, 2018a, b). Appendix A gives a brief description of the formulation of the spectral problem for stationary, cylindrical equilibria in terms of a second order ODE, which forms the foundation for the ROC code ánd for the analytical results of Goedbloed & Keppens (2022). Key to our further discussion of the SARI are the MHD continua, which are singular frequency ranges that organise the spectrum. Because of the rotational equilibrium flow, all frequencies are Doppler shifted according to the Doppler range,
@@
Note that, because $v_{z}=0$ in our set-up, the Doppler shift vanishes for axisymmetric ($m=0$) perturbations. The four singular ranges are dubbed forward and backward Alfvén and slow continuum, as the frequencies lie in a band on the real axis. The continua are given by,
@@
which are the Doppler shifted Alfvén and slow continua. From hereon, $\Omega_{A,S}^{\pm}$ denotes the closed range of continuum frequencies expressed by $\{\Omega_{A,S}^{\pm}(r)\}$ in Goedbloed & Keppens (2022). In accretion disks with subthermal magnetic fields, these Alfvén and slow frequency ranges almost coincide because of incompressibility (if $\gamma\rightarrow+\infty$, they exactly coincide). Our numerical results do always account for compressibility effects, as we adopt a ratio of specific heats $\gamma=5/3$ for the disk plasma. Whereas ±$\pm\omega_{A}$ are always separated for $m=0$ modes like the MRI, the continua can overlap for non-axisymmetric modes. Exactly this phenomenon is responsible for many of the interesting properties of SARIs. For large values of $m$, $\Omega_{0}$ dominates over $\omega_{A}$ and the forward and backward continua overlap to a large degree.
@@ 3.1 Accretion disk instabilities
The MHD spectrum of a typical weakly magnetised, near-Keplerian disk for axisymmetric $m=0$ modes contains the MRI, among other modes. The structure of such a spectrum is well-known (Keppens et al., 2002; Blokland et al., 2005; Das et al., 2018; Goedbloed & Keppens, 2022), and is shown in Fig. 1 (a) for a weak field with equal poloidal and toroidal components and $k=70$. Vertically, we find a finite sequence of overstable discrete MRI modes (and their damped counterparts), and horizontally, oscillating Alfvénic/slow modes clustering to either edges of the Alfvén continua, indicated by red and black segments on the real axis. In fact, the oscillating modes clustering to the edges ±$\pm\omega_{A2}$ belong in the sequence of MRI modes when listed according to the increasing number of zeroes of their eigenfunctions but they are stable (in a local MRI analog, the radial wavenumber $q$ becomes larger so as to stabilise the MRI). Two sequences of fast modes cluster towards infinity, but they are not shown in this figure. The up-down symmetry of the spectrum is a standard feature of ideal MHD where all operators are self-adjoint (Goedbloed et al., 2019), but the damped MRI variant is of course of limited physical significance for accretion disk turbulence (but it is significant in the initial value problem). The spectrum is, however, not left-right symmetric as a result of background flow. In particular, the axisymmetric MRI is not exponentially unstable (note the slight deviation from the imaginary axis), in contrast with local WKB predictions (Balbus & Hawley, 1991), although there exists a local dispersion relation that does reproduce this behaviour (Blokland et al., 2005). A typical eigenfunction, here the real part of $v_{r}$ for the most unstable eigenvalue, is shown in the inset (the imaginary part looks similar). The perturbation is localised at the inner boundary and has no zeroes inside the domain. Moving down through the sequence of MRIs, the number of zeroes increases and the perturbations become global over the entire domain. They are hence sensitive to the boundary conditions.
@@
Although a complete spectroscopic picture of non-axisymmetric accretion disk instabilities had been missing until recently, it has been known for some time that a wavenumber $m\neq 0$ produces an additional branch of discrete unstable modes if the Alfvén continua overlap. These modes resonate with both Alfvén continua and corotate with the disk at the local Doppler frequency (Matsumoto & Tajima, 1995; Ogilvie & Pringle, 1996; Curry & Pudritz, 1996). Since a weakly magnetised, thin disk rotates super-Alfvénically ($v_{A1}\ll 1\approx\Omega_{1}$), Goedbloed & Keppens (2022) termed them super-Alfvénic rotational instabilities in their general analysis, where they showed that the two branches consist of infinitely many discrete modes clustering towards the internal edges of the Doppler-shifted continua $\Omega_{A}^{+}$ and $\Omega_{A}^{-}$. These modes are all unstable, in contrast with the finite number of unstable MRIs on Fig. 1 (a). Figure 1 (b) features a typical SARI spectrum for $m=-10$, considering a smaller domain with $\delta=0.1$ and the same equilibrium as for Panel (a). Note the large Doppler shift that moves the oscillation frequencies to negative values, making the continua $\Omega_{A}^{\pm}$ (red and black) overlap in this super-Alfvénic regime. For these counter-rotating SARIs (since $m<0$), both $\Omega_{0}$ and $\Omega_{A}^{\pm}$ are decreasing, and the inner SARI eigenfrequencies cluster to $\Omega_{A}^{-}(r_{1})$ whereas the outer SARIs cluster to $\Omega_{A}^{+}(r_{2})$ (and vice-versa for co-rotating modes with $m>0$). The most unstable SARIs are indicated in the figure, and we note that the inner SARI growth rate is largest – a general feature of disks with outwardly-decreasing equilibrium profiles (for an explanation, we refer to Sect. 5.4). The fact that both branches have oscillation frequencies $\sigma$ inside the Doppler range (gray) implies that these modes corotate with (or against) the local flow at the Doppler corotation radius $r_{*}$ where $\Omega_{0}(r_{*})=\sigma$. Since $\sigma$ is also in one of the continua $\Omega_{A}^{\pm}$, each mode has one Alfvénic resonance, which has direct bearing on the eigenfunctions shown in the two insets of Panel (b), with vertical lines denoting the resonances: the inner $v_{r}$ eigenfunction (orange) is localised between the inner boundary and the resonance with the backward Alfvén continuum; the outer eigenfunction (green) between the resonance with the forward Alfvén continuum and the outer boundary. Both peak at the location of resonance with the local Doppler frequency. Clearly, the names ‘inner’ and ‘outer’ also reflect the localisation of the eigenfunctions (Ogilvie & Pringle, 1996). Skin currents at the resonant location shield the mode from one boundary, acting as a virtual wall (Goedbloed & Keppens, 2022). The number of zeroes of the eigenfunctions (‘radial wavenumber’) again increases down the sequence, but it is less straightforward to number them accordingly (Goedbloed & Keppens, 2022).
@@ 3.2 Radially extended disks
Even though the radial domain in Fig. 1 (b) was chosen to be small, the SARI eigenfunctions are not global because they are limited by an Alfvén resonance. In some sense, each mode actually has a second Alfvén resonance outside of the domain at a fictitious radius if the continua are extended. Intuitively, these additional resonances are encapsulated when the domain is enlarged, and the result might be hybrid modes that have properties of both inner and outer SARIs: they are localised between the two resonances, and hence shielded from the boundaries. It is clear that for a real accretion disk, these boundaries are unphysical. Physically relevant modes should hence preferentially be insensitive to these artificial constraints, like the modes in radially extended disks proposed here.
In reality, the matter is more complicated, but indeed, overlapping forward and backward shifted Alfvén continua can produce modes with two resonances, forming an unprecedented 2D region in the complex eigenfrequency plane of overstable quasi-modes, as recently reported by Goedbloed & Keppens (2022). In regimes with a large overlap between $\Omega_{A}^{\pm}$ (if $\delta$ is not small, or $k$ and $m$ large), a wealth of modes appears above the Doppler range. These modes have a non-zero complementary energy $W_{\text{com}}$, a complex-valued quantity that should vanish for true normal modes (Goedbloed, 2018a) (see Appendix A). This complementary energy quantifies the addition of energy required to excite these modes; they are almost eigenoscillations like normal modes, but not exactly. However, for these modes, $W_{\text{com}}$ is very small, sometimes even below machine precision, rendering the resulting near-eigenmodes numerically indistinguishable from actual normal modes – hence Goedbloed & Keppens (2022) termed them quasi-modes. By allowing a small but finite $W_{\text{com}}$, dense regions in the complex plane containing such modes appear, from hereon named continua of quasi-modes or quasi-continua. In practice, non-zero $W_{\text{com}}$ implies that the total pressure perturbation shows a minute jump at the location where a shooting method links up two numerical solutions starting from the left and right boundaries, but this is numerically virtually unnoticeable at small values of $W_{\text{com}}=10^{-12}$ that we have here. Physically, $W_{\text{com}}$ is the additional energy required to bring the mode into resonance with the time dependency $\exp(-i\omega t)$. Such tiny excitation energies as for the quasi-modes in this work could easily be provided by ambient energy fluctuations in turbulent disks, so that there is no observable difference with a normal mode. Moreover, in actual non-linear numerical MHD simulations, one has always round-off and truncation errors, related to the spatio-temporal discretisations used. Deviations from machine precision can be sufficient to introduce non-zero complementary energy at any time during the run.
@@ 4 Quasi-continuum modes
The above discussion on quasi-modes and their small but finite complementary energy represents an unusual numerical challenge to a finite-element-based solver like Legolas. Because of the Galerkin decomposition of eigenmodes, the boundary conditions and eigenfunction continuity (even differentiability) will be satisfied per construction, and will not show a minute jump associated with non-zero $W_{\text{com}}$. If a quasi-mode is found, it is a priori truly indistinguishable from any other mode in the spectrum. However, a change in resolution will always slightly shift the quasi-mode eigenvalues found previously, whereas convergence towards a solution generally requires that an eigenvalue/eigenvector pair changes minimally when increasing the resolution. This regular convergence behaviour is indeed true for discrete modes like MRI or the inner and outer, discrete SARI variants.
@@
Despite these difficulties, it is possible to validate the quasi-continuum results from Goedbloed & Keppens (2022) with Legolas with the combined effort of multiple solvers implemented in the code. Figure 2 overlays some of our results on Fig. 15 of Goedbloed & Keppens (2022) for the same equilibrium. Panel (a) shows the $m=1$ co-propagating SARI spectrum as found using the QR-invert algorithm for increasing resolution $n_{\text{grid}}\in\{150,200,300,500\}$. The spectra are overplotted on the $W_{\text{com}}\leq 10^{-12}$ (black) and $10^{-8}$ regions calculated with the ROC code and show a good match: the QR-invert algorithm locates the left and right edges of the quasi-continuum, as well as additionally finding the discrete inner and outer SARI modes clustering to the continua and disappearing in a sea of quasi-modes. For increasing grid resolution, the upper edge of the quasi-continuum as calculated by Legolas remains more or less stationary, but the lower edge shifts downwards, thereby increasingly resolving the lateral edges. Since the independent Spectral Web approach can identify the entire black shaded region as having a minute complementary energy, we must conclude that the lower edge of the quasi-continuum obtained by the QR algorithm is hence an artificial feature of the limited resolution. The correspondence between the quasi-continuum edges and the black region for $W_{\text{com}}=10^{-12}$ (small compared to the normalised eigenfunctions) indicates that Legolas operates on a similar accuracy. It should be noted that low resolution versions of Fig. 1 (b) also show a quasi-continuum like structure between the two SARI branches, but it becomes smaller as the resolution is increased and its upper edge approaches the real axis. On the other hand, the upper edges of true quasi-continua quickly converge to their final values when the resolution is increased. We use this observation to define the first criterion for the detection of quasi-continua with Legolas: a QR-solver locates edges of the quasi-continuum, and with increasing resolution, the upper edge converges whereas the lateral edges are increasingly resolved.
@@
To show that the regions delimited by the QR-invert solver actually contain infinitely many unstable modes, we use the Arnoldi shift-invert solver (ARPACK, Lehoucq et al. 1998), which finds a predefined number of eigenvalues closest to some complex target frequency $\omega_{t}$. We define a grid of target shift frequencies that sample the 2D region of interest, where we use the quasi-continuum edges found by the QR algorithm. Thanks to the recent upgrades to the Legolas code (Claes & Keppens, 2023), this can now easily be done at high radial resolutions: we choose a resolution of $n=1000$ gridpoints for the equilibria and look for 100 eigenvalues in the neighbourhood of each shift. Figure. 2 (b) shows the result of several of such runs for $m=1$. An annulus of eigenvalues is found around each shift in the quasi-continuum, with the inner and outer radius depending on the location, resolution and number of eigenvalues. However, it is clear that the annuli are only artificial: the eigenvalue clouds found by nearby shifts overlap, allowing for the 2D quasi-continua to be filled completely (as we do in Fig. 3). This observation can be used as a second criterion for defining quasi-continua with Legolas: when applying shift-invert within the bounds determined by the QR-algorithm, arbitrarily many modes can be located in the direct neighbourhood of the shift.
@@
Panels (c) and (d) of Fig. 2 show three quasi-mode eigenfunctions, sampled at various locations in the quasi-continuum as indicated in Panel (b). All three eigenfunctions $\Re(v_{r})$ resemble wave packages with a different number of oscillations and shifted to various locations in the disk. Note how the number of oscillations increases as $\Im(\omega)$ decreases, but the eigenfunctions remain radially localised – in contrast to the MRI, which becomes more global as the number of oscillations increases. Like the discrete SARI eigenfunctions, each quasi-mode eigenfunction peaks at the local Doppler corotation radius (dashed, gray) and has one or two resonances with the Alfvén continua (red and black, dotted). The eigenmode in Panel (c) has resonances with both continua. At these locations, large skin currents appear that effectively shield the mode from the radial boundaries (Goedbloed & Keppens, 2022). This eigenmode then represents a truly radially localised wave package moving at the local Doppler speed, insensitive to the boundaries. How local a quasi-continuum mode is, is determined by the separation of these two resonances: if the overlap of $\Omega_{A}^{\pm}$ is large, as it is for high-$m$ modes, the separation is very small and modes are hence radially very localised (but still global in nature). The two quasi-modes in Panel (d): the orange and green eigenfunctions only resonate with the forward (red) and backward (black) Alfvén continuum, respectively. In this sense, modes like these are the quasi-continuum equivalent of discrete SARI modes, where the second Alfvén resonance can now either be inside the domain or not. An inspection of the other eigenfunctions (not shown) reveals that the Lagrangian perturbations along and perpendicular to the magnetic field are largest, highlighting the slow/Alfvénic nature of the instability.

### v2

@@ 2.2 Eigenvalue problem
corresponding to a Fourier decomposition in the spatial and temporal domains with complex Fourier coefficients $\hat{f}(r)$, real azimuthal and vertical wavenumbers $m$ and $k$, respectively, and complex eigenfrequencies $\omega=\sigma+i\nu$. Due to the $\theta$-periodicity of the cylindrical equilibrium, $m$ is quantised to integer values, whereas the vertical wavenumber $k$ can take arbitrary values in this case. The set of eigenoscillations of this equilibrium is fully determined with the wavenumbers $m$ and $k$.
We calculate all MHD eigenspectra in this work using the Legolas [^1] https://legolas.science. The code version used in this work is open-source under the GPL-3.0 license and available on GitHub at https://github.com/n-claes/legolas/tree/v2.0.0. code developed by Claes et al. (2020) and De Jonghe et al. (2022) and recently upgraded in terms of solver and memory performance (Claes & Keppens, 2023). It uses a finite-element approach to solve the following generalised matrix eigenvalue problem obtained from the linearised MHD equations, where the Fourier coefficients become eigenfunctions:
@@
Here, $\mathbf{X}$ is an 8-component eigenvector before its Galerkin decomposition (dropping the hat on the Fourier coefficients), containing eigenfunctions $(\rho,v_{r},v_{\theta},v_{z},T,a_{r},a_{\theta},a_{z})$, with $\rho$ the density perturbation, $T$ the plasma temperature perturbation and $\mathbf{a}$ the magnetic vector potential, which fixes the perturbed magnetic field $\mathbf{B}=\nabla\times\mathbf{a}$ and enforces the monopole constraint $\nabla\cdot\mathbf{B}$ exactly. Legolas constructs the matrices A and B from the equilibrium variations, the chosen finite-element base functions in the eigenmode representations, and the required physical effects like viscosity, resistivity, Hall MHD, etc. (De Jonghe et al., 2022). As this work is limited to ideal MHD, the matrices A and B are modified from the static, linear MHD case with only flow and gravity as additional effects. The matrix system is then solved using various available solvers depending on the requirements. In this work, we use the QR-invert algorithm to produce complete eigenspectra, and the Arnoldi shift-invert method to search for eigenvalues in the neighbourhood of some complex number (Claes & Keppens, 2023). The mathematical formulation of the eigenvalue problem is closed by fixing the boundary conditions at radius $1$ and $1+\delta$. By default, Legolas assumes perfectly conducting, solid walls as boundaries, requiring $v_{r}=0$ and $a_{\theta}=a_{z}=0$ at those locations. As explained in Goedbloed & Keppens (2022), these solid-wall boundaries will affect the detailed variation of MRIs, while SARIs are fully agnostic of these boundary conditions, an important distinction between both mode categories.
To substantiate our Legolas results, we combine them in several figures with results obtained with the complementary ROC code, which employs a shooting method (Goedbloed, 2018a, b). Appendix A gives a brief description of the formulation of the spectral problem for stationary, cylindrical equilibria in terms of a second order ODE, which forms the foundation for the ROC code ánd for the analytical results of Goedbloed & Keppens (2022). Key to our further discussion of the SARI are the MHD continua, which are singular frequency ranges that organise the spectrum. Because of the rotational equilibrium flow, all frequencies are Doppler-shifted according to the Doppler range:
@@
Note that, because $v_{z}=0$ in our set-up, the Doppler shift vanishes for axisymmetric ($m=0$) perturbations. The four singular ranges are dubbed forward and backward Alfvén and slow continua, as the frequencies lie in a band on the real axis. The continua are given by,
@@
which are the Doppler-shifted Alfvén and slow continua. From here on, $\Omega_{A,S}^{\pm}$ denotes the closed range of continuum frequencies expressed by $\{\Omega_{A,S}^{\pm}(r)\}$ in Goedbloed & Keppens (2022). In accretion disks with subthermal magnetic fields, these Alfvén and slow frequency ranges almost coincide because of incompressibility (if $\gamma\rightarrow+\infty$, they exactly coincide). Our numerical results do always account for compressibility effects, as we adopt a ratio of specific heats $\gamma=5/3$ for the disk plasma. Whereas ±$\pm\omega_{A}$ are always separated for $m=0$ modes like the MRI, the continua can overlap for non-axisymmetric modes. Exactly this phenomenon is responsible for many of the interesting properties of SARIs. For large values of $m$, $\Omega_{0}$ dominates over $\omega_{A}$ and the forward and backward continua overlap to a large degree.
@@ 3.1 Accretion disk instabilities
The MHD spectrum of a typical weakly magnetised, near-Keplerian disk for axisymmetric $m=0$ modes contains the MRI, among other modes. The structure of such a spectrum is well-known (Keppens et al., 2002; Blokland et al., 2005; Das et al., 2018; Goedbloed & Keppens, 2022), and is shown in Fig. 1 (a) for a weak field with equal poloidal and toroidal components and $k=70$. Vertically, we find a finite sequence of overstable discrete MRI modes (and their damped counterparts), and horizontally, we find oscillating Alfvénic/slow modes clustering to either edge of the Alfvén continua, indicated by red and black segments on the real axis. In fact, the oscillating modes clustering to the edges ±$\pm\omega_{A2}$ belong to the sequence of MRI modes when listed according to the increasing number of zeroes of their eigenfunctions but they are stable (in a local MRI analog, the radial wavenumber $q$ becomes larger, stabilising the MRI). Two sequences of fast modes cluster towards infinity, but they are not shown in the figure. The up-down symmetry of the spectrum is a standard feature of ideal MHD where all operators are self-adjoint (Goedbloed et al., 2019), but the damped MRI variant is of course of limited physical significance for accretion disk turbulence (but it is significant in the initial value problem). The spectrum is, however, not left-right symmetric as a result of background flow. In particular, the axisymmetric MRI is not exponentially unstable (note the slight deviation from the imaginary axis), in contrast with local WKB predictions (Balbus & Hawley, 1991), although there exists a local dispersion relation that does reproduce this behaviour (Blokland et al., 2005). A typical eigenfunction, here the real part of $v_{r}$ for the most unstable eigenvalue, is shown in the inset (the imaginary part looks similar). The perturbation is localised at the inner boundary and has no zeroes inside the domain. Moving down through the sequence of MRIs, the number of zeroes increases and the perturbations become global over the entire domain. They are hence sensitive to the boundary conditions.
@@
Although a complete spectroscopic picture of non-axisymmetric accretion disk instabilities has been missing until recently, it has been known for some time that a wavenumber $m\neq 0$ produces an additional branch of discrete unstable modes if the Alfvén continua overlap. These modes resonate with both Alfvén continua and corotate with the disk at the local Doppler frequency (Matsumoto & Tajima, 1995; Ogilvie & Pringle, 1996; Curry & Pudritz, 1996). Since a weakly magnetised, thin disk rotates super-Alfvénically ($v_{A1}\ll 1\approx\Omega_{1}$), Goedbloed & Keppens (2022) termed them super-Alfvénic rotational instabilities in their general analysis, where they showed that the two branches consist of infinitely many discrete modes clustering towards the internal edges of the Doppler-shifted continua $\Omega_{A}^{+}$ and $\Omega_{A}^{-}$. These modes are all unstable, in contrast with the finite number of unstable MRIs in Fig. 1 (a). Figure 1 (b) features a typical SARI spectrum for $m=-10$, considering a smaller domain with $\delta=0.1$ and the same equilibrium as that in Panel (a). Note the large Doppler shift that moves the oscillation frequencies to negative values, making the continua $\Omega_{A}^{\pm}$ (red and black) overlap in this super-Alfvénic regime. For these counter-rotating SARIs (since $m<0$), both $\Omega_{0}$ and $\Omega_{A}^{\pm}$ are decreasing, and the inner SARI eigenfrequencies cluster to $\Omega_{A}^{-}(r_{1})$ whereas the outer SARIs cluster to $\Omega_{A}^{+}(r_{2})$ (and vice-versa for co-rotating modes with $m>0$). The most unstable SARIs are indicated in the figure, and we note that the inner SARI growth rate is largest – a general feature of disks with outwardly-decreasing equilibrium profiles (for an explanation, we refer to Sect. 5.4). The fact that both branches have oscillation frequencies $\sigma$ inside the Doppler range (gray) implies that these modes corotate with (or against) the local flow at the Doppler corotation radius $r_{*}$ where $\Omega_{0}(r_{*})=\sigma$. Since $\sigma$ is also in one of the continua $\Omega_{A}^{\pm}$, each mode has one Alfvénic resonance, which has direct bearing on the eigenfunctions shown in the two insets of Panel (b), with vertical lines denoting the resonances: the inner $v_{r}$ eigenfunction (orange) is localised between the inner boundary and the resonance with the backward Alfvén continuum, the outer eigenfunction (green) between the resonance with the forward Alfvén continuum and the outer boundary. Both peak at the location of resonance with the local Doppler frequency. Clearly, the names ‘inner’ and ‘outer’ also reflect the localisation of the eigenfunctions (Ogilvie & Pringle, 1996). Skin currents at the resonant location shield the mode from one boundary, acting as a virtual wall (Goedbloed & Keppens, 2022). The number of zeroes of the eigenfunctions (the ‘radial wavenumber’) again increases down the sequence, but it is less straightforward to number them accordingly (Goedbloed & Keppens, 2022).
@@ 3.2 Radially extended disks
Even though the radial domain in Fig. 1 (b) is chosen to be small, the SARI eigenfunctions are not global because they are limited by an Alfvén resonance. In some sense, each mode actually has a second Alfvén resonance outside of the domain at a fictitious radius if the continua are extended. Intuitively, these additional resonances are encapsulated when the domain is enlarged, and the result might be hybrid modes that have properties of both inner and outer SARIs: they are localised between the two resonances, and hence shielded from the boundaries. It is clear that for a real accretion disk, these boundaries are unphysical. Physically relevant modes should hence preferentially be insensitive to these artificial constraints, like the modes in radially extended disks proposed here.
In reality, the matter is more complicated, but indeed, overlapping forward and backward shifted Alfvén continua can produce modes with two resonances, forming an unprecedented 2D region in the complex eigenfrequency plane of overstable quasi-modes, as recently reported by Goedbloed & Keppens (2022). In regimes with a large overlap between $\Omega_{A}^{\pm}$ (if $\delta$ is not small, or $k$ and $m$ are large), a wealth of modes appears above the Doppler range. These modes have a non-zero complementary energy $W_{\text{com}}$, a complex-valued quantity that should vanish for true normal modes (Goedbloed, 2018a) (see Appendix A). This complementary energy quantifies the addition of energy required to excite these modes; they are almost eigenoscillations like normal modes, but not exactly. However, for these modes, $W_{\text{com}}$ is very small, sometimes even below machine precision, rendering the resulting near-eigenmodes numerically indistinguishable from actual normal modes – hence Goedbloed & Keppens (2022) termed them quasi-modes. By allowing a small but finite $W_{\text{com}}$, dense regions in the complex plane containing such modes appear, from here on named continua of quasi-modes or quasi-continua. In practice, non-zero $W_{\text{com}}$ implies that the total pressure perturbation shows a minute jump at the location where a shooting method links up two numerical solutions starting from the left and right boundaries, but this is numerically virtually unnoticeable at the small values of $W_{\text{com}}=10^{-12}$ that we have here. Physically, $W_{\text{com}}$ is the additional energy required to bring the mode into resonance with the time dependency $\exp(-i\omega t)$. Such tiny excitation energies as for the quasi-modes in this work could easily be provided by ambient energy fluctuations in turbulent disks, so that there is no observable difference from a normal mode. Moreover, in actual non-linear numerical MHD simulations, one always has round-off and truncation errors, related to the spatio-temporal discretisations used. Deviations from machine precision can be sufficient to introduce non-zero complementary energy at any time during the run.
@@ 4 Quasi-continuum modes
The above discussion on quasi-modes and their small but finite complementary energy represents an unusual numerical challenge to a finite-element-based solver like Legolas. Because of the Galerkin decomposition of eigenmodes, the boundary conditions and eigenfunction continuity (even differentiability) will be satisfied per construction, and will not show a minute jump associated with non-zero $W_{\text{com}}$. If a quasi-mode is found, it is a priori truly indistinguishable from any other mode in the spectrum. However, a change in resolution will always slightly shift the quasi-mode eigenvalues found previously, whereas convergence towards a solution generally requires that an eigenvalue/eigenvector pair change minimally when increasing the resolution. This regular convergence behaviour is indeed true for discrete modes like MRI or the inner and outer, discrete SARI variants.
@@
Despite these difficulties, it is possible to validate the quasi-continuum results from Goedbloed & Keppens (2022) with Legolas with the combined effort of multiple solvers implemented in the code. Figure 2 overlays some of our results in Fig. 15 of Goedbloed & Keppens (2022) for the same equilibrium. Panel (a) shows the $m=1$ co-propagating SARI spectrum found using the QR-invert algorithm for increasing resolution $n_{\text{grid}}\in\{150,200,300,500\}$. The spectra are overplotted on the $W_{\text{com}}\leq 10^{-12}$ (black) and $10^{-8}$ regions calculated with the ROC code and show a good match: the QR-invert algorithm locates the left and right edges of the quasi-continuum, as well as additionally finding the discrete inner and outer SARI modes clustering to the continua and disappearing in a sea of quasi-modes. For increasing grid resolution, the upper edge of the quasi-continuum as calculated by Legolas remains more or less stationary, but the lower edge shifts downwards, thereby increasingly resolving the lateral edges. Since the independent Spectral Web approach can identify the entire black shaded region as having a minute complementary energy, we must conclude that the lower edge of the quasi-continuum obtained by the QR algorithm is hence an artificial feature of the limited resolution. The correspondence between the quasi-continuum edges and the black region for $W_{\text{com}}=10^{-12}$ (small compared to the normalised eigenfunctions) indicates that Legolas operates on a similar accuracy. It should be noted that low-resolution versions of Fig. 1 (b) also show a quasi-continuum like structure between the two SARI branches, but it becomes smaller as the resolution is increased and its upper edge approaches the real axis. On the other hand, the upper edges of true quasi-continua quickly converge to their final values when the resolution is increased. We use this observation to define the first criterion for the detection of quasi-continua with Legolas: a QR-solver locates edges of the quasi-continuum, and with increasing resolution, the upper edge converges whereas the lateral edges are increasingly resolved.
@@
To show that the regions delimited by the QR-invert solver actually contain infinitely many unstable modes, we use the Arnoldi shift-invert solver (ARPACK, Lehoucq et al. 1998), which finds a predefined number of eigenvalues closest to some complex target frequency $\omega_{t}$. We define a grid of target shift frequencies that sample the 2D region of interest, where we use the quasi-continuum edges found by the QR algorithm. Thanks to the recent upgrades to the Legolas code (Claes & Keppens, 2023), this can now easily be done at high radial resolutions: we choose a resolution of $n=1000$ gridpoints for the equilibria and look for 100 eigenvalues in the neighbourhood of each shift. Figure. 2 (b) shows the result of several of such runs for $m=1$. An annulus of eigenvalues is found around each shift in the quasi-continuum, with the inner and outer radii depending on the location, resolution, and number of eigenvalues. However, it is clear that the annuli are only artificial: the eigenvalue clouds found by nearby shifts overlap, allowing for the 2D quasi-continua to be filled completely (as we do in Fig. 3). This observation can be used as a second criterion for defining quasi-continua with Legolas: when applying shift-invert within the bounds determined by the QR-algorithm, arbitrarily many modes can be located in the direct neighbourhood of the shift.
@@
Panels (c) and (d) of Fig. 2 show three quasi-mode eigenfunctions, sampled at various locations in the quasi-continuum as indicated in Panel (b). All three eigenfunctions $\Re(v_{r})$ resemble wave packages having a different number of oscillations and shifted to various locations in the disk. Note how the number of oscillations increases as $\Im(\omega)$ decreases, but the eigenfunctions remain radially localised – in contrast to the MRI, which becomes more global as the number of oscillations increases. Like the discrete SARI eigenfunctions, each quasi-mode eigenfunction peaks at the local Doppler corotation radius (dashed, gray) and has one or two resonances with the Alfvén continua (red and black, dotted). The eigenmode in Panel (c) has resonances with both continua. At these locations, large skin currents appear that effectively shield the mode from the radial boundaries (Goedbloed & Keppens, 2022). This eigenmode then represents a truly radially localised wave package moving at the local Doppler speed, insensitive to the boundaries. How local a quasi-continuum mode is, is determined by the separation of these two resonances: if the overlap of $\Omega_{A}^{\pm}$ is large, as it is for high-$m$ modes, the separation is very small and the modes are hence radially very localised (but still global in nature). As for the two quasi-modes in Panel (d), the orange and green eigenfunctions only resonate with the forward (red) and backward (black) Alfvén continua, respectively. In this sense, modes like these are the quasi-continuum equivalent of discrete SARI modes, where the second Alfvén resonance can now be either inside the domain or not. An inspection of the other eigenfunctions (not shown) reveals that the Lagrangian perturbations along and perpendicular to the magnetic field are the largest, highlighting the slow/Alfvénic nature of the instability.